# Approximating evidence via bounded harmonic means

### v1

@@ 4.0.1 Example 1: Multivariate Gaussian distributions
In this symmetric and unimodal example, Mixture THAMES yields poorer results in terms of precision compared to the other estimators, as shown in Figure 2. This is likely due to the additional Monte Carlo step required by Mixture THAMES to estimate the intersection volume between the ellipsoid and the HPD region. Furthermore, THAMES is comparable to ECMLE despite the former using an optimal level set and the latter a local model-free approximation. Both THAMES and Mix-THAMES were implemented using the optimal configurations recommended by Metodiev et al. [12, 11]. In THAMES, the ellipsoid radius was set to $r=\sqrt{d+1}$, and Mixture THAMES used the truncation level determined by minimizing the Kolmogorov distance between the truncated, standardized negative log-posterior and the $\chi^{2}_{d}$ distribution.
Figure 1: (Example 1) Comparison of log-marginal-likelihood estimates across different HPD levels using ECMLE. Each box plot was created by performing 100 replications of the algorithms each of which used $10^{5}$ posterior draws. The red dashed line represents the exact value of the marginal likelihood.
Figure 2: (Example 1) Comparison of log-marginal-likelihood estimates obtained using three different estimators. The dashed red line indicates the exact marginal likelihood value, and the number of MCMC draws is $T=10^{5}$ for each of the 100 replications of the algorithms, with $\alpha=0.75$.
4.0.2 Example 2: Mixture of multivariate Gaussian distributions
@@
The exact marginal likelihood is also available
@@
In this multimodal toy example, Figure 3 illustrates how the HPD simulations are contributing to the approximate HPD region. In the case of ECMLE, the HPD regions are covered by a union of two ellipsoids and therefore ECMLE is using the target topology. Predictably, THAMES fails to account for bimodality and consequently includes some extremely low posterior density regions. The truncated version of THAMES (Mixture THAMES) manages to handle this issue since, by construction, the mixture based proposal allows elimination of these low posterior density regions while requiring an independent Monte Carlo evaluation of its volume.
@@
Figure 4: (Example 2) Comparison of marginal-likelihood estimates across different HPD levels using ECMLE method, for a mixture of two Gaussian posteriors. The red dashed line indicates the exact log-marginal-likelihood value.
@@
To reinforce this initial evaluation, we estimated directly the variance of ECMLE and the Mixture THAMES. Since both estimators are unbiased, the differences between these estimators were contained in the square expectation
@@
$\mathbb{E}\left[\hat{Z}^{-2}\right]=\frac{1}{TV(A)^{2}}\int_{A}\frac{1}{\pi(\theta)L(\theta)}\text{d}\theta\,$, (6)
@@
that can be approximated by a Monte Carlo estimate based on a Uniform sample on $A$. Figure 5 confirms the above conclusions and shows that a value of $\alpha$ in the vicinity of $80\%$ achieves better precision than at other HPD levels. ECMLE also exhibits less variability or variance in estimates compared to THAMES versions, see Figure 6.
@@
Figure 5: (Example 2) Comparison of the variance of the estimators based on the formula (6) and approximated by MCMC for $8$ HPD levels, from $\alpha=0.1$ to $\alpha=0.99$. The number of MCMC draws is $10^{5}$.
Figure 6: (Example 2) Comparison of log-marginal-likelihood estimates for each estimator, illustrated using the Multivariate Mixture of Gaussians example. Here, ECMLE method outperforms the other methods by providing the most stable and accurate estimates. For all 100 replications of the algorithms, $T=10^{5}$ MCMC draws are used, and the HPD level is set to $\alpha=0.75$.
@@
For a similar setting involving data generated from a mixture of $K$ Gaussian distributions as a prior, Figure 7 shows the coverage of the parameter posterior samples using ECMLE, Mixture THAMES and THAMES ellipsoid for $K=4$, $K=6$, or $K=8$ mixture components. Note that each panel involves a single data sample of $n=20$ points. For $\alpha=75\%$, Figure 7’s top left side shows four red circles indicating the $75\%$ high-density region that has been perfectly covered by ECMLE. The blue points represent the posterior sample points within the $\alpha$-HPD which are properly located inside ECMLE ellipsoids, while gray ones fall outside these regions. Figure 7’s middle and right panels display THAMES and Mixture THAMES ellipsoids by red lines. In the right panel case, among the points within the ellipsoid, blue points enjoy one of the $\alpha$% highest posterior density values. By comparing each plot with the left-hand one, we observe that some of the $\alpha\%$-HPD points have not been covered by Mixture THAMES. Furthermore, in the middle panel, a significant part of non HPD posterior samples fall in the ellipsoid highlighting the numerous cases where the ellipsoid does not adequately cover the high-density region. Further, ECMLE ($K=6$) covered $71.82\%$ of the $\alpha$-HPD region, Mixture THAMES ($K=6$) covered $67.83\%$ and the THAMES ellipsoid involved $82.51\%$ of the posterior samples.
@@
Moreover, as shown by Figure 7, ECMLE achieves accurate HPD approximation through its adaptive boundary-aware design. Unlike fixed geometric shapes, ECMLE places ellipsoids exclusively at HPD sample locations and calculates each semi-axes to ensure maximal coverage within high-density regions while precisely respecting the HPD boundaries. This adaptive semi-axes calculation allows the method to conform locally to arbitrary posterior geometries (whether multimodal, skewed, or irregularly shaped) without the geometric constraints of THAMES that may systematically include low-density areas or miss complex boundary structures.
@@
Figure 7: (Example 2) Coverage of the posterior samples: (Left) ECMLE; (Middle) THAMES; (Right) Mixture THAMES; for three Gaussian mixture posterior distributions with $K=4$, $K=6$, and $K=8$ components. The Blue points represent posterior draws considered by each method for estimating the marginal likelihood, and gray points fall outside when the HPD level is 75% HPD region. Each panel corresponds to a dataset of size $n=20$ observations and $T=5\times 10^{4}$ iid simulations from the posterior. All methods are based on the same MCMC simulations and dataset.
We extended our analysis to a higher dimensional experiment where $d=10$. As shown in Figure 8, ECMLE method continued to perform slightly better, showing the lowest variance and bias. While Mixture THAMES performed well, it required a significant increase in computation, needing $10^{5}$ additional simulations to calculate intersection volumes for this dimension.
Figure 8: (Example 2) (Right) Comparison of log-evidence estimations obtained for a mixture of Gaussians in dimension 10, with a zoom (Left) to compare ECMLE and Mixture THAMES.
4.0.3 Example 3: Rosenbrock Distribution
@@
In this example, we examine the performances of the different methods previously mentioned for a posterior distribution with a boomerang shape called the Rosenbrock distribution that has often been used as a benchmark in the Bayesian computational literature [6, 23, 15]. In the general case, each entry $X_{i}\in\mathbb{R}^{d-1}$ ($i=1,\dots,n$) follows the density
@@
where $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{d})^{\top}\in\mathbb{R}^{d}$ is the parameter vector, and $a,b\in\mathbb{R}^{d-1}$ are fixed constants. This formulation captures a chain of quadratic dependencies, starting from $\theta_{1}$ without a direct observation and propagating through subsequent dimensions.
The likelihood for $n$ independent observations is then
@@
To ensure a proper posterior, we assign priors as follows: $\theta_{1}\sim\mathcal{N}(0,\sigma_{1}^{2})$ (weakly informative) and $\pi(\theta_{j})=1$ for $j=2,\dots,d$ (flat priors). This setup yields a well-defined joint posterior distribution.
In our specific example, we leverage sufficient statistics to simplify the model. Let $\bar{Y}=(\bar{Y}_{1},\bar{Y}_{2},\dots,\bar{Y}_{d})^{\top}\in\mathbb{R}^{d}$ denote the vector of the sample means of $n$ observations, where $\bar{Y}_{j}=\frac{1}{n}\sum_{i=1}^{n}x_{i,j}$ for $j=1,\dots,d-1$ (noting that the original observations are in $\mathbb{R}^{d-1}$, but we extend to include a direct term for completeness in this variant). These sample means serve as sufficient statistics for the parameters under the Gaussian likelihood, condensing the data without loss of information.
@@
This structure maintains the Rosenbrock dependencies while incorporating observations across all dimensions, making $\bar{Y}$ a natural choice for inference. The full likelihood density becomes
@@
Here, we adopt improper flat priors on all parameters: $\pi(\boldsymbol{\theta})\propto 1$. This choice simplifies calculations but requires caution, as it can lead to improper posteriors in some cases; however, the added observation dimension ensures propriety in this setup.
@@
$Z=\int p(\bar{Y}\mid\boldsymbol{\theta})\,\pi(\boldsymbol{\theta})\,d\boldsymbol{\theta}=1$
@@
(see Supplement Exact Evidence Computation of Rosenbrock Example for the detailed calculation). This constant evidence highlights a key property of the model under improper priors: the marginal likelihood is data-independent, which can be useful for certain theoretical analyses but may not hold in more constrained settings.
@@
Figure 9 illustrates the different covering of the $\alpha$-HPD region by the three methods and highlights the difficulties of THAMES in fitting the non-linear structure of the posterior surface. In such a case, since the covariance matrix of the observations does not convey useful information about the shape of the posterior, Mixture THAMES is also missing a part of the HPD region. In contrast, ECMLE extends further and better represents the full structure of the HPD region, including its separate branches. This results in a higher precision for ECMLE as illustrated by Figure 10, while the original THAMES fails to capture the true value of the evidence.
@@
Figure 9: (Example 3) Approximations of the $\alpha=0.75$ HPD region for the Rosenbrock posterior distribution.
Figure 10: (Example 3) Comparison of the marginal likelihood estimates obtained using ECMLE and Mixture THAMES for three data dimensions of the Rosenbrock distribution. The dashed red line indicates the exact marginal likelihood value. Each method is based on $10^{5}$ MCMC draws over 100 replications. The data sample size is $n=20$ for $d=2$, and $n=200$ for $d=5$ and $d=10$.
Finally, the execution time of each method has been monitored in order to determine the respective computational efficiencies. For each example, we performed 100 repetitions of MCMC draws ($T=10^{5}$), which were provided to all algorithms. In each repetition, the same underlying dataset and posterior sample were used. Table 1 summarizes the root mean squared error (RMSE) and the product of the RMSE and the algorithm execution time. The RMSE quantifies how far the estimates deviate from the true value of the evidence; however, it should not be overemphasized, as none of the algorithms are explicitly designed to minimize this measure of efficiency.
Table 1: Method comparisons for different examples
Method
Example Metric THAMES Mix THAMES ECMLE
Multivariate Gaussian RMSE 0.00327 0.0116 0.00333
RMSE $\times$ Time 0.0016 0.6154 0.0020
Mixture of Multivariate RMSE 0.0130 0.0056 0.0038
Gaussian RMSE $\times$ Time 0.0064 0.4204 0.0037
Rosenbrock (2D) RMSE 0.1375 0.0135 0.0051
RMSE $\times$ Time 0.0820 0.1720 0.0081
Rosenbrock (5D) RMSE 23.4308 0.0508 0.0405
RMSE $\times$ Time 14.5145 0.7384 0.4430
Rosenbrock (10D) RMSE 1441.8007 0.3977 0.1596
RMSE $\times$ Time 1008.3400 5.4055 4.7213
From Table 1, we observe that THAMES generally provides the fastest computation times, especially in low-dimensional cases, but its RMSE increases significantly for higher-dimensional, multimodal, and more challenging examples like the Rosenbrock function. Mixture THAMES reduces RMSE considerably compared to the original THAMES but at a much higher computational cost. ECMLE consistently achieves the lowest RMSE in most examples, particularly in higher dimensions, while maintaining a reasonable runtime. Overall, when considering the RMSE $\times$ Time metric as a measure of efficiency, ECMLE tends to provide the best trade-off between accuracy and computational cost across all tested scenarios.

### v2

@@ 4 Numerical Illustrations
In this symmetric and unimodal example, tTHAMES yields poorer results in terms of precision compared to the other estimators, as shown in Figure 2. This is likely due to the additional Monte Carlo step required by tTHAMES to estimate the intersection volume between the ellipsoid and the HPD region. Furthermore, THAMES is comparable to ECMLE despite the former using an optimal level set and the latter a local model-free approximation. Both THAMES and tTHAMES were implemented using the optimal configurations recommended by Metodiev et al. [13, 12]. In THAMES, the ellipsoid radius was set to $r=\sqrt{d+1}$, and tTHAMES used the truncation level determined by minimizing the Kolmogorov distance between the truncated, standardized negative log-posterior and the $\chi^{2}_{d}$ distribution. We used the authors’ original code for both THAMES and tTHAMES. For tTHAMES, we compute the estimator over the set $A^{\prime}_{\hat{\theta},\hat{\Sigma},r,\alpha}$ (Equation 6), excluding the label-switching corrections for mixture models. Notably, PWK performs well in this setting, benefiting from the low dimensionality and the symmetric, near-circular shape of the posterior, which aligns naturally with its spherical shell partitioning scheme.
Figure 1: (Example 1) Comparison of log-marginal-likelihood estimates across different HPD levels using ECMLE. Each box plot was created by performing 100 replications of the algorithms, each using $10^{5}$ posterior draws for the same observed data. The red dashed line represents the exact value of the marginal likelihood.
Figure 2: (Example 1) Evidence ratio $\widehat{Z}/Z$ for the bivariate Gaussian model with $n\in\{20,50,100\}$. Each boxplot is based on 100 datasets. To equalize computation time across methods, the number of posterior draws was set to: THAMES (850,000), tTHAMES (38,000), ECMLE (425,000), and ePWK (48,000).
Example 2: Mixture of multivariate Gaussian distributions
@@
The exact marginal likelihood is available as
@@
In this multimodal toy example, Figure 3 illustrates how the HPD simulations are contributing to the approximate HPD region. In the case of ECMLE, the HPD regions are covered by a union of two ellipsoids and therefore ECMLE is using the target topology. Predictably, THAMES fails to account for bimodality and consequently includes some extremely low posterior density regions. The truncated version of THAMES manages to handle this issue since, by construction, the mixture based proposal allows elimination of these low posterior density regions while requiring an independent Monte Carlo evaluation of its volume.
@@
Figure 4: (Example 2) Comparison of marginal-likelihood estimates across different HPD levels using ECMLE method, for a mixture of two Gaussian posteriors. Each boxplot was created by performing 100 replications of the algorithm, each using $10^{5}$ posterior draws for the same observed data. The red dashed line indicates the exact log-marginal-likelihood value.
Figure 4 shows the sensitivity of ECMLE to the HPD level $\alpha$ in the bimodal setting, confirming that values around $0.75$ yield the most stable estimates.
@@
To reinforce this initial evaluation, we estimated directly the variance of ECMLE and tTHAMES. Since both estimators are unbiased, the differences between these estimators were contained in the square expectation
@@
$\mathbb{E}\left[\hat{Z}^{-2}\right]=\frac{1}{TV(\mathcal{R})^{2}}\int_{\mathcal{R}}\frac{1}{\pi(\boldsymbol{\theta})L(\boldsymbol{\theta})}\,\mathrm{d}\boldsymbol{\theta}\,$, (7)
@@
where $\mathcal{R}$ denotes the instrumental region ($\mathcal{E}$ for ECMLE and $A$ for tTHAMES). This quantity can be approximated by a Monte Carlo estimate based on a uniform sample on $\mathcal{R}$. Figure 5 confirms the above conclusions and shows that a value of $\alpha$ in the vicinity of $80\%$ achieves better precision than at other HPD levels. ECMLE also exhibits less variability in estimates compared to the other methods across all sample sizes (n = 20, 50, 100), as shown in Figure 6, confirming its stability in multimodal settings.
@@
Figure 5: (Example 2) Comparison of the variance of the estimators based on the formula (7), approximated by Monte Carlo for different HPD levels, from $\alpha=0.1$ to $\alpha=0.99$. The number of posterior draws is $10^{5}$.
Figure 6: (Example 2) Boxplots of the evidence ratio ($\widehat{Z}/Z$) for the 2D Gaussian mixture model with $K=2$ and $n\in\{20,50,100\}$, based on $M=100$ datasets. The ideal value is 11. To ensure a fair comparison at matched runtime, we use per-dataset posterior draws: THAMES (850,000), tTHAMES (25,500), ECMLE (420,000), and ePWK (25,000).
@@
Extending to priors with $K\in\{4,6,8\}$ Gaussian mixture components, with observations drawn from $\mathcal{N}(0,\Sigma)$, Figure 7 shows the coverage of the posterior samples using ECMLE, tTHAMES, and THAMES. Each panel corresponds to a single dataset of $n=20$ observations. For $\alpha=0.75$, Figure 7, ECMLE with $K=4$ (top left), shows four red ellipsoids indicating the $75\%$ high-density region that has been perfectly covered by ECMLE. The blue points represent posterior samples within the $\alpha$-HPD that fall inside the ECMLE ellipsoids, while gray ones fall outside these regions. Figure 7, THAMES and tTHAMES (middle and right columns), display their ellipsoids by red lines. For tTHAMES (right column), among the points within the ellipsoid, blue points correspond to those with the $\alpha$ highest posterior density values. By comparing with ECMLE (left column), we observe that some of the $\alpha$-HPD points have not been covered by tTHAMES. Furthermore, for THAMES (middle column), a significant portion of non-HPD posterior samples fall within the ellipsoid, highlighting cases where a single ellipsoid fails to adequately capture the high-density region. For $K=6$, ECMLE covered $71.82\%$ of the $\alpha$-HPD region, tTHAMES covered $67.83\%$, and the THAMES ellipsoid included $82.51\%$ of all posterior samples.
@@
Moreover, as shown by Figure 7, ECMLE achieves accurate HPD approximation through its adaptive boundary-aware design. Unlike fixed geometric shapes, ECMLE places ellipsoids exclusively at HPD sample locations and calculates each semi-axis to ensure maximal coverage within high-density regions while precisely respecting the HPD boundaries. This adaptive semi-axes calculation allows the method to conform locally to arbitrary posterior geometries (whether multimodal, skewed, or irregularly shaped) without the geometric constraints of THAMES that may systematically include low-density areas or miss complex boundary structures.
@@
Figure 7: (Example 2) Coverage of the posterior samples: (Left) ECMLE; (Middle) THAMES; (Right) tTHAMES; for three Gaussian mixture posterior distributions with $K=4$, $K=6$, and $K=8$ components. The blue points represent posterior draws considered by each method for estimating the marginal likelihood, and gray points fall outside when the HPD level is 75% HPD region. Each panel corresponds to a dataset of size $n=20$ observations and $T=5\times 10^{4}$ iid simulations from the posterior. All methods are based on the same MCMC simulations and dataset.
To further evaluate the performance of the estimators across varying model complexity and sample sizes, Figure 8 presents a comprehensive comparison of the evidence ratio ($\widehat{Z}/Z$) for THAMES, tTHAMES, ECMLE, and ePWK across nine configurations of the Gaussian mixture model, with $K\in\{4,6,8\}$ mixture components and sample sizes $n\in\{20,50,100\}$. All methods were calibrated to operate under approximately the same computational budget, ensuring a fair comparison based on estimation accuracy rather than runtime differences. Across all configurations, ECMLE consistently achieves estimates closest to the true evidence value with the lowest variability. As the number of mixture components increases, the advantage of ECMLE becomes more pronounced, reflecting its ability to adapt to increasingly complex multimodal posterior geometries. In contrast, THAMES exhibits substantial bias and high variance, particularly for larger $K$, as its single ellipsoid fails to capture the multiple modes of the posterior. The tTHAMES improves upon THAMES but still shows greater variability than ECMLE, especially for smaller sample sizes. The ePWK estimator, while competitive in lower-dimensional symmetric cases, struggles as the posterior complexity grows, highlighting its sensitivity to deviations from ideal geometric conditions.
Figure 8: (Example 2) Comparison of evidence ratio ($\widehat{Z}/Z$) for THAMES, tTHAMES, ECMLE, and ePWK across nine configurations of the Gaussian mixture model: $K\in\{4,6,8\}$ mixture components and sample sizes $n\in\{20,50,100\}$. All methods were calibrated so that each runs under approximately the same computational budget, ensuring that comparisons reflect estimation accuracy rather than differences in runtime.
Example 3: Rosenbrock Distribution
@@
In this example, we examine the performances of the different methods previously mentioned for a posterior distribution with a boomerang shape called the Rosenbrock distribution that has often been used as a benchmark in the Bayesian computational literature ([6, 24, 16]).
More precisely, each entry $X_{i}\in\mathbb{R}^{d-1}$ ($i=1,\dots,n$) follows the density
@@
where $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{d})^{\top}\in\mathbb{R}^{d}$ is the parameter vector, and $a,b\in\mathbb{R}^{d-1}$ are fixed constants. This formulation captures a chain of quadratic dependencies, starting from $\theta_{1}$ without a direct observation and propagating through subsequent dimensions. The likelihood for $n$ independent observations is then
@@
In our specific example, we leverage sufficient statistics to simplify the model. We consider a $d$-dimensional formulation where each parameter $\theta_{j}$ has a corresponding observation $\bar{Y}_{j}$, enabling both exact posterior sampling and a closed-form marginal likelihood. Let $\bar{Y}=(\bar{Y}_{1},\bar{Y}_{2},\dots,\bar{Y}_{d})^{\top}\in\mathbb{R}^{d}$ denote the vector of sample means, where $\bar{Y}_{j}=\frac{1}{n}\sum_{i=1}^{n}x_{i,j}$ for $j=1,\dots,d$. Here $\bar{Y}_{1}$ directly informs $\theta_{1}$, while the remaining observations $\bar{Y}_{2},\ldots,\bar{Y}_{d}$ retain the Rosenbrock nonlinear dependence on preceding parameters.
@@
This structure maintains the Rosenbrock dependencies while incorporating observations across all dimensions. The full likelihood density becomes
@@
Here, we adopt improper flat priors on all parameters: $\pi(\boldsymbol{\theta})\propto 1$. Although improper priors may lead to improper posteriors, here the posterior is proper since each parameter $\theta_{j}$ is directly informed by its corresponding observation $\bar{Y}_{j}$.
@@
$Z=\int p(\bar{Y}\mid\boldsymbol{\theta})\,\pi(\boldsymbol{\theta})\,\text{d}\boldsymbol{\theta}=1$
@@
(see supplementary material for the detailed calculation). This constant evidence highlights a key property of the model under improper priors: the marginal likelihood is data-independent, which can be useful for certain theoretical analyses but may not hold in more constrained settings.
@@
Figure 9 illustrates the different coverages of the $\alpha$-HPD region by the three methods and highlights the difficulties of THAMES in fitting the non-linear structure of the posterior surface. In such a case, since the covariance matrix of the observations does not convey useful information about the shape of the posterior, tTHAMES is also missing a part of the HPD region. In contrast, ECMLE extends further and better represents the full structure of the HPD region, including its separate branches.
@@
Figure 9: (Example 3) Approximations of the $\alpha=0.75$ HPD region for the two-dimensional Rosenbrock posterior distribution.
Figure 10 presents a comprehensive evaluation of marginal likelihood estimates for the Rosenbrock distribution across dimensions $d\in\{2,5,10\}$ and sample sizes $n\in\{20,50,100\}$, based on 100 independently generated datasets. Due to their extremely high variances in this challenging setting, THAMES and ePWK are excluded from this comparison; their results are provided in supplementary material for completeness. All methods were calibrated to operate under comparable computational budgets, with method-specific numbers of posterior draws adjusted accordingly.
In dimensions $d=5$ and $d=10$, the tTHAMES method encounters difficulties when computing the intersection volume using simple Monte Carlo approximation: the uniform samples drawn within the ellipsoid often fail to fall inside the HPD-truncated intersection region, resulting in a computed volume of zero. To address this, the method employs a multi-step strategy, shifting the ellipsoid center from the posterior mean to the posterior mode and shrinking the radius of the initial ellipsoid. While this adaptive approach allows the algorithm to proceed, it highlights the challenging geometry of the Rosenbrock distribution in higher dimensions and the limitations of relying on a single global ellipsoid to capture such complex posterior structures.
Across all configurations, ECMLE consistently provides the most accurate and stable estimates, with values tightly concentrated around the true marginal likelihood ($Z=1$). This demonstrates the effectiveness of ECMLE’s adaptive ellipsoid covering in capturing the curved, banana-shaped geometry of the Rosenbrock distribution.
Figure 10: (Example 3) Boxplots of marginal likelihood estimates for the Rosenbrock distribution posterior in dimensions $d\in\{2,5,10\}$ and sample sizes $n\in\{20,50,100\}$, based on $M=100$ independently generated datasets. For each dataset, the remaining methods use the same posterior sample, but with method-specific numbers of draws chosen so that all methods operate under a comparable computational budget. For $d=2$: tTHAMES (25,500), ECMLE (420,000). For $d=5$: tTHAMES (48,000), ECMLE (350,000). For $d=10$: tTHAMES (67,000), ECMLE (28,000). The dashed horizontal line marks the exact marginal likelihood ($Z=1$).